# Entanglement generation and scaling from noisy quenches across a quantum critical point

### v2

@@ IV Entanglement generation from noisy ramped quenches
Let us now turn our attention to the effect from noise on entanglement generation when quenching across a critical point of the TFI chain. As discussed in Sec. I, the experimentally relevant protocol is to repeat the quench many times, then taking an average over the outcome, in effect averaging over the noise distribution. Computationally, a “brute-force” option is to repeatedly solve the stochastic Schrödinger equation for a large number of single noise realizations, then forming an average [66]. A simpler, more elegant way is to solve an exact noise master equation [67, 44, 45, 68] for the noise-averaged density matrix $\rho_{k}(t)=\langle\rho_{\eta,k}(t)\rangle$, with $\rho_{\eta,k}(t)$ the density matrix for a given mode $k$ and white-noise realization $\eta(t)$, Eq. (10). Explicitly,
$\dot{\rho}_{k}(t)=-i[H_{0,k}(t),\rho_{k}(t)]-\frac{\xi^{2}}{2}[\sigma^{z},[\sigma^{z},\rho_{k}(t)]].$
@@
The noise-averaged nearest-neighbor and next-nearest-neighbor concurrence $\langle C_{\l ,\l +1}\rangle$ and $\langle C_{\l ,\l +2}\rangle$, respectively, are displayed in Fig. 2 as a function of $h_{0}(t)$ for $\tau=10$ and three different values of noise intensity $\xi$ and noiseless case $\xi=0$ corresponding to Fig. 1. With the exception of a very weak intensity, $\xi=0.005$ in Fig. 2(a), the noise is seen to reduce the entanglement for almost all values of $h_{0}(t)$. This is altogether expected since noise generically acts as a source of decoherence.
To estimate the amount of entanglement reduction due to noise, we have plotted the noise-averaged concurrence in Fig. 3 as a function of $\tau$ for four different noise intensities $\xi$, for quenches from $h_{i}=-30$ to $h_{f}=5$ $h_{f}=30]$ for $\langle C_{\l ,\l +1}\rangle$ $\langle C_{\l ,\l +2}\rangle]$. As shown above, the concurrence between spins separated by an odd number of lattice spacings get entangled for finite $h_{f}$. To make this feature visible in the presence of noise, we have chosen a smaller value of $h_{f}$ in panel 3(a) (as compared to panel 3(b)) since $\langle C_{\l ,\l +1}\rangle$ vanishes for larger values. The result displayed in Fig. 3 is qualitatively anticipated: The larger the noise intensity, the larger the entanglement reduction.
@@
Figure 4: (Color online) (a) Power law scaling $\tau_{c}(\xi)\sim\xi^{-\delta}$ with exponent $\delta=0.666\pm 0.001\approx 2/3$, where $\xi$ is the noise intensity and $\tau_{c}(\xi)$ is the critical time scale at which the noise-averaged concurrence $\langle C_{\l ,\l +2}\rangle$ vanishes. The inset shows the linear scaling of the maximum next-nearest-neighbor concurrence with $\xi^{2}$. (b) Noise-averaged concurrence between next-nearest-neighbor spins as functions of $\ln(\tau)$ for a ramped quench with added white noise for different values of noise intensity $\xi$, with quench interval from $h_{i}=-30$ to $h_{f}=30$.
It is interesting to contrast our result for $\langle C_{\l ,\l +2}\rangle$ with the noiseless case discussed by Sengupta and Sen [33], where concurrence between spins separated by an even number of lattice bonds is analyzed. Whereas the noiseless concurrence $C_{\l ,\l +2}$ in Fig. 3(b) vanishes when the time scale of the ramped quench goes to infinity, in agreement with Ref. [33], the noise-averaged concurrence $\langle C_{\l ,\l +2}\rangle$ vanishes already at finite values of the time scale. Importantly, note that the entanglement reduction for fixed noise intensity increases with increasing $\tau$, indicating that for larger values of $\tau$ there is more time for noise to become effective and, by that, cause disentanglement. An analysis reveals that noise with intensity $\xi\geq 0.05$ wipes out $\langle C_{\l ,\l +2}\rangle$ for all $\tau$ (cf. Fig. 2(b) when $\tau=10$), while for $0<\xi<0.05$ there is a critical time scale, call it $\tau_{c}(\xi)$, for which the spins remain entangled when $\tau_{0}<\tau<\tau_{c}(\xi)$ (with $\tau_{0}\approx 1.96$ defined above).
@@
with $\delta=0.666\pm 0.001\approx 2/3$. Comparison with $\tau_{c}(\xi)$ for $\xi=0.004,0.006$ and $0.01$ as seen in Fig. 3(b), implies that $\tau_{c}(\xi)\approx 0.12\,(\xi^{2})^{-2/3}.$ We have chosen to present the scaling in Eq. (15) using the squared noise intensity $\xi^{2}$ that determines the strength of the white-noise correlations, Eq. (10). This brings out an intriguing connection with the anti-Kibble-Zurek behavior as discussed in Ref. [50]: These authors found that the optimal quench time that minimizes the production of defects scales with $\xi^{2}$ as a power law with exponent $\delta=2/3$, the same exponent as in our Eq. (15). While our quench protocol is different from that in Ref. [50] - with the quench crossing both critical points in the TFI chain instead of just one - the very same protocol as ours, applied to the XY chain (of which the TIFM is a special case), was later found to also yield the same exponent $\delta=2/3$ [51]. Whether the identity of the exponents, both found in quenches that emulate standard quantum annealing schemes, is only coincidental or signals a hidden link between the two phenomena - minimization of KZM defects and vanishing of next-nearest neighbor concurrence in the presence of noise - remains an open question.
Making a log plot (not shown) of the red curve for the noiseless case in Fig. 3(b), we immediately verify the KZM scaling law $C_{\l ,\l +2}\sim 1/\sqrt{\tau}$ established in [33] for slow quenches, $\tau\gtrsim 150$. As displayed in Fig. 4(b), the curves from Fig. 3(b) with nonzero noise instead scale logarithmically with $\tau$ before vanishing at $\tau_{c}(\xi)$,
@@
The result in Eq. (16) differs from the anti-Kibble-Zurek scaling of defect formation first found in Ref. [50] and obtained for the same noisy quenches [51] as explored by us. The authors of Refs. [50, 51] showed that noise contributes a term $\sim\tau$, added to the usual noiseless KZM scaling term $\sim 1/\sqrt{\tau}$. Whereas our power-law scaling for the vanishing of the next-nearest-neighbor concurrence in the presence of noise, Eq. (15), may hint at a link to defect formation, the difference of the logarithmic scaling in Eq. (16) from that in Refs. [50, 51] appears to rule out that the two phenomena are quantitatively connected.
It is important to note that the oscillations in $\langle C_{l,l+1}\rangle$ for finite values of $h_{f}$ (cf. Fig. 3(a)) preclude a direct scaling analysis similar to that for $\langle C_{l,l+2}\rangle$. This is the reason why we do not present scaling results for $\langle C_{l,l+1}\rangle$.
@@ V Summary and outlook
Collecting our results for slow noisy quenches across the critical points of the TFI chain, we are led to infer that the direct link between KZM defect formation and entanglement generation suggested by the results in Refs. [14, 19, 33, 34, 35] gets broken by noise. Specifically, the known KZM scaling $\sim 1/\sqrt{\tau}$ of the next-nearest neighbor concurrence [33], with $\tau$ the time scale of the quench, does not apply when noise is present. Instead, we find a logarithmic scaling $\sim\ln(1/\tau)$. Additional evidence that entanglement generation is nontrivially connected - if at all connected - to the defect formation comes from a comparison with results by Dutta et al. [50] and Gao at al. [51]. These authors found that the scaling of the defect density is governed by two independent terms: one noisy term $\sim\tau$, and one KZM term $\sim 1/\sqrt{\tau}$, again different from the logarithm uncovered by our work.
@@ Acknowledgement
Appendix
V.1 Defect generation under noisy ramp fields
As discussed in Sec. 1, defect generation under noisy ramp fields exhibits an anti-Kibble-Zurek (AKZ) behavior when crossing a quantum critical point, whereby slower driving results in a higher density of defects [50, 51, 52, 53, 54, 55, 56].
@@ V.1 Defect generation under noisy ramp fields
As the ramp time scale increases, noise-induced defects dominate the nonadiabatic dynamics, leading to a growth of the defect density with $\tau$. This is indicative of the AKZ regime, where enhancing the ramp time scale results in a higher density of defects in the system. In the limit of a very large values of $\tau$, $n_{\xi}$ is completely governed by the AKZ contribution. In summary, the defect generation is controlled by two competing mechanisms: (i) non-adiabatic defect density, suppressed by increasing the time scale $\tau$ of the ramp; and (ii) noise-induced defect density, amplified by increasing $\tau$. As shown in Refs. [50, 51], the interplay of these two competing influences yields a minimum of the density of defects ($n^{min}_{\xi}$) at an optimum annealing time $\tau_{opt}$ that scales as $\tau_{opt}\propto\tau^{-\delta}$, with $\delta=0.666\pm 0.001\simeq 2/3$.
@@
V.2 Computing concurrence for time-evolved states of the TFI chain
@@
$C_{\l ,m}(t)=\mbox{max}(0,\lambda_{1}(t)\!-\!\lambda_{2}(t)\!-\!\lambda_{3}(t)\!-\!\lambda_{4}(t)).$ (A1)
@@
$R_{\l ,m}\!=\!\sqrt{\sqrt{\rho_{\l ,m}(t)}\tilde{\rho}_{\l ,m}(t)\sqrt{\rho_{\l ,m}(t)}},$ (A2)
@@
Figure 5: (Color online) The logarithm of the density of defects $n_{\xi}$ versus time scale $\tau$ of a ramped quench crossing a quantum critical point in the transverse field Ising chain for different noise intensities $\xi$. The behavior of the density of defects for different values of $\tau$ and $\xi$ illustrates anti-Kibble-Zurek behavior in the presence of a noisy ramp field.
@@
[…]
@@
$\kappa_{44}$ $=$ $-\langle s^{z}_{l}\rangle+\langle s^{z}_{l}s^{z}_{m}\rangle+\frac{1}{4},$ (A8)
@@
Singling out a reference spin for notational transparency, say at site $\l =1$, and using the JW transformation in Eq. (II), the two-point functions in Eq. (V.2) can be expressed in terms of multipoint correlators of the JW fermions,
@@
$\langle s_{1}^{x}s_{1+r}^{y}\rangle$ $\!=\!$ $-\frac{i}{4}\langle B_{1}A_{2}B_{2}...A_{r}B_{r}B_{r+1}\rangle,$ (A9)
@@
$A_{j}=c_{j}^{\dagger}+c_{j},\ \ B_{j}=c_{j}^{\dagger}-c_{j},\ \ j=1,\ldots,r+1.$ (A10)
As for the one-point function $\langle s^{z}_{l}\rangle$ in Eq. (V.2), one has
$\langle s^{z}_{l}\rangle=\frac{1}{2}\langle B_{l}A_{l}\rangle.$ (A11)
The expectation values in Eq. (V.2) are defined with respect to the time-evolved fermion states that stand in for the spin states after the JW transformation. For example,
$\langle A_{1}B_{2}\ldots\rangle=\mbox{Tr}(\rho(t)A_{1}B_{2}\ldots)$ (A12)
with $\rho(t)=\otimes_{k}\rho_{k}(t)$, where $\rho_{k}(t)$ is the density operator for mode $k$ in a pure [mixed] ensemble of noise-free [noise-averaged] states from ramped quenches; cf. text before [Eq. (III)] [Eq. (IV)]. Note that in the case of a pure ensemble, we used the notation $\rho_{k}(t)=\rho_{0,k}(t)$ in Sec. III. Also note that the product over density matrices for single modes runs over $k=\pm(2i-1)\pi/N,i=1,2,\ldots,N/2$; cf. text after Eq. (2).
The transverse spin correlation functions in Eq. (V.2) can be written on the generic form
$\langle s_{1}^{\alpha}s_{1+r}^{\beta}\rangle=D_{r}^{\alpha\beta}\langle{{\phi_{1}}{\phi_{2}}{\phi_{3}}...\phi_{2r-2}\phi_{2r-1}{\phi_{2r}}}\rangle,$ (A13)
with $D_{r}^{xx}=\frac{1}{4},D_{r}^{yy}=(-1)^{r}\frac{1}{4},D_{r}^{xy}=-\frac{i}{4},$ and $D_{r}^{yx}=(-1)^{r}\frac{i}{4}$, where each operator $\phi_{j},j=1,2,\ldots,2r$, is identified with either $A_{j}$ or $B_{j}$ according to the corresponding expression in Eq. (V.2). Note further that $\langle s_{1}^{z}s_{1+r}^{z}\rangle$ has the simple form
$\langle s_{1}^{z}s_{1+r}^{z}\rangle=\frac{1}{4}\langle{{\phi_{1}}{\phi_{2}}{\phi_{3}}{\phi_{4}}}\rangle,$ (A14)
@@
The expressions in Eqs. (A13) and (A14) are expedient as they allow us to express the multipoint correlators in Eq (V.2) in terms of Pfaffians. This scheme was first exploited by Barouch and McCoy [73] in their study of the XY chain, using a result by Caianiello and Fubini [76] obtained via Wick’s theorem.
@@
where we have written the skew-symmetric matrices on standard abbreviated form. By combining Eqs. (V.2), (A13), (A21), and (A26), replacing the choice l=1l\!=\!1 by the collective index $l$, it is straightforward to show that the correlation functions for nearest-neighbour spins, r=1r\!=\!1, and next-nearest-neighbour spins, r=2r\!=\!2, are given by
@@
$\langle s^{x}_{l}s^{y}_{l+1}\rangle$ $=$ $-\frac{i}{4}\langle B_{l}B_{l+1}\rangle,$ (A27)
@@
$\langle s^{y}_{l}s^{y}_{l+2}\rangle$ $=$ $\frac{1}{4}\large[\langle B_{l+1}A_{l}\rangle\langle B_{l+2}A_{l+1}\rangle\!-\!\langle A_{l}A_{l+1}\rangle\langle B_{l+1}B_{l+2}\rangle-\langle B_{l+2}A_{l}\rangle\langle B_{l+2}A_{l+2}\rangle\large],$ (A28)
@@
$\langle B_{l}B_{\l +r}\rangle$ $=$ $\langle c_{l}^{\dagger}c_{\l +r}^{\dagger}\rangle+\langle c_{l}c_{\l +r}\rangle-\langle c_{l}^{\dagger}c_{\l +r}\rangle-\langle c_{l}c_{\l +r}^{\dagger}\rangle,$ (A29)
@@
where, by way of translational invariance, $r=0,\pm 1,\pm 2$ cover all cases in Eqs. (V.2) and (V.2). We obtain closed expressions for the spin correlations in these equations by calculating the fermion two-point functions in Eq. (V.2). Explicitly, let us consider the two-point function $\langle c_{\l }^{\dagger}c_{\l +r}^{\dagger}\rangle$, suppressing time arguments for notational simplicity, and write
$\langle c_{\l }^{\dagger}c_{\l +r}^{\dagger}\rangle=\mbox{Tr}(\rho\,c_{\l }^{\dagger}c_{\l +r}^{\dagger}),$ (A30)
where, like in Eq. (A12), $\rho=\otimes_{k}\,\rho_{k}$ with $\rho_{k}$ the density operator for mode $k$. By Fourier transforming, one obtains from Eq. (A30),
$\langle c^{\dagger}_{\ell}c^{\dagger}_{\ell+r}\rangle$ $=$ $\frac{1}{N}\sum_{p,q}e^{-i(p\ell+q(\ell+r))}\langle c^{\dagger}_{p}c^{\dagger}_{q}\rangle$
$=$ $\frac{1}{N}\sum_{p,q}e^{-i(p\ell+q(\ell+r))}\mbox{Tr}(\rho\,c^{\dagger}_{p}c^{\dagger}_{q})$

### v3

@@ IV Entanglement generation from noisy ramped quenches
Let us now turn our attention to the effect from noise on entanglement generation when quenching across a critical point of the TFI chain. As discussed in Sec. I, the experimentally relevant protocol is to repeat the quench many times, then taking an average over the outcome, in effect averaging over the noise distribution. Computationally, a “brute-force” option is to repeatedly solve the stochastic Schrödinger equation for a large number of single noise realizations, then forming an average [66]. A simpler, more elegant way is to solve an exact noise master equation [67, 44, 45, 68] for the noise-averaged density matrix $\rho_{k}(t)=\langle\rho_{\eta,k}(t)\rangle$, with $\rho_{\eta,k}(t)$ the density matrix for a given mode $k$ and white-noise realization $\eta(t)$ [Eq. (10)]. Explicitly,
$\dot{\rho}_{k}(t)=-i[H_{0,k}(t),\rho_{k}(t)]-\frac{\xi^{2}}{2}[\sigma^{z},[\sigma^{z},\rho_{k}(t)]],$
@@
The noise-averaged nearest-neighbor and next-nearest-neighbor concurrence $\langle C_{\l ,\l +1}\rangle$ and $\langle C_{\l ,\l +2}\rangle$, respectively, are displayed in Fig. 2 as a function of $h_{0}(t)$ for $\tau=10$ and three different values of noise intensity $\xi$ and noiseless case $\xi=0$ corresponding to Fig. 1. With the exception of a very weak intensity, $\xi=0.005$ in Fig. 2(a), the noise is seen to reduce the entanglement for almost all values of $h_{0}(t)$. This is expected since noise generically acts as a source of decoherence.
To estimate the amount of entanglement reduction due to noise, we have plotted the noise-averaged concurrence in Fig. 3 as a function of $\tau$ for four different noise intensities $\xi$, for quenches from $h_{i}=-30$ to $h_{f}=5$ $h_{f}=30]$ for $\langle C_{\l ,\l +1}\rangle$ $\langle C_{\l ,\l +2}\rangle]$. As shown above, the concurrence between spins separated by an odd number of lattice spacings get entangled for finite $h_{f}$. To make this feature visible in the presence of noise, we have chosen a smaller value of $h_{f}$ in Fig. 3(a) [as compared to Fig. 3(b)] since $\langle C_{\l ,\l +1}\rangle$ vanishes for larger values. The result displayed in Fig. 3 is qualitatively anticipated: The larger the noise intensity, the larger the entanglement reduction.
@@
Figure 4: (Color online) (a) Power-law scaling $\tau_{c}(\xi)\sim\xi^{-\delta}$ with exponent $\delta=0.666\pm 0.001\approx 2/3$, where $\xi$ is the noise intensity and $\tau_{c}(\xi)$ is the critical time scale at which the noise-averaged concurrence $\langle C_{\l ,\l +2}\rangle$ vanishes. The inset shows the linear scaling of the maximum next-nearest-neighbor concurrence with $\xi^{2}$. (b) Noise-averaged concurrence between next-nearest-neighbor spins as functions of $\ln(\tau)$ for a ramped quench with added white noise for different values of noise intensity $\xi$, with quench interval from $h_{i}=-30$ to $h_{f}=30$.
It is interesting to contrast our result for $\langle C_{\l ,\l +2}\rangle$ with the noiseless case discussed by Sengupta and Sen [33], where concurrence between spins separated by an even number of lattice bonds is analyzed. Whereas the noiseless concurrence $C_{\l ,\l +2}$ in Fig. 3(b) vanishes when the time scale of the ramped quench goes to infinity, in agreement with Ref. [33], the noise-averaged concurrence $\langle C_{\l ,\l +2}\rangle$ vanishes already at finite values of the time scale. Importantly, note that the entanglement reduction for fixed noise intensity increases with increasing $\tau$, indicating that for larger values of $\tau$ there is more time for noise to become effective and, by that, cause disentanglement. An analysis reveals that noise with intensity $\xi\geq 0.05$ wipes out $\langle C_{\l ,\l +2}\rangle$ for all $\tau$ [cf. Fig. 2(b) when $\tau=10$], while for $0<\xi<0.05$ there is a critical time scale, call it $\tau_{c}(\xi)$, for which the spins remain entangled when $\tau_{0}<\tau<\tau_{c}(\xi)$ (with $\tau_{0}\approx 1.96$ defined above).
@@
with $\delta=0.666\pm 0.001\approx 2/3$. Comparison with $\tau_{c}(\xi)$ for $\xi=0.004,0.006$, and $0.01$ as seen in Fig. 3(b), implies that $\tau_{c}(\xi)\approx 0.12\,(\xi^{2})^{-2/3}.$ We have chosen to present the scaling in Eq. (15) using the squared noise intensity $\xi^{2}$ that determines the strength of the white-noise correlations [Eq. (10)]. This brings out an intriguing connection with the anti-Kibble-Zurek behavior as discussed in Ref. [50]: These authors found that the optimal quench time that minimizes the production of defects scales with $1/\xi^{2}$ as a power law with exponent $\delta=2/3$, the same exponent as in our Eq. (15). While our quench protocol is different from that in Ref. [50] - with the quench crossing both critical points in the TFI chain instead of just one - the very same protocol as ours, applied to the XY chain (of which the TFI chain is a special case), was later found to also yield the same exponent $\delta=2/3$ [51]. Whether the identity of the exponents, both found in quenches that emulate standard quantum annealing schemes, is only coincidental or signals a hidden link between the two phenomena - minimization of KZM defects and vanishing of next-nearest neighbor concurrence in the presence of noise - remains an open question.
Making a log plot (not shown) of the red curve for the noiseless case in Fig. 3(b), we immediately verify the KZM scaling law $C_{\l ,\l +2}\sim 1/\sqrt{\tau}$ established in Ref. [33] for slow quenches, $\tau\gtrsim 150$. As displayed in Fig. 4(b), the curves from Fig. 3(b) with nonzero noise instead scale logarithmically with $\tau$ before vanishing at $\tau_{c}(\xi)$,
@@
The result in Eq. (16) differs from the anti-Kibble-Zurek scaling of defect formation first found in Ref. [50] and obtained for the same noisy quenches [51] as explored by us. The authors of Refs. [50, 51] showed that noise contributes a term $\sim\tau$, added to the usual noiseless KZM scaling term $\sim 1/\sqrt{\tau}$. Whereas our power-law scaling for the vanishing of the next-nearest-neighbor concurrence in the presence of noise [Eq. (15)] may hint at a link to defect formation, the difference of the logarithmic scaling in Eq. (16) from that in Refs. [50, 51] appears to rule out that the two phenomena are quantitatively connected.
It is important to note that the oscillations in $\langle C_{l,l+1}\rangle$ for finite values of $h_{f}$ [cf. Fig. 3(a)] preclude a direct scaling analysis similar to that for $\langle C_{l,l+2}\rangle$. This is the reason why we do not present scaling results for $\langle C_{l,l+1}\rangle$.
@@ V Summary and outlook
Collecting our results for slow noisy quenches across the critical points of the TFI chain, we are led to infer that the direct link between KZM defect formation and entanglement generation suggested by the results in Refs. [14, 19, 33, 34, 35] gets broken by noise. Specifically, the known KZM scaling $\sim 1/\sqrt{\tau}$ of the next-nearest neighbor concurrence [33], with $\tau$ the time scale of the quench, does not apply when noise is present. Instead, we find a logarithmic scaling $\sim\ln(1/\tau)$. Additional evidence that entanglement generation is nontrivially connected - if at all connected - to the defect formation comes from a comparison with the AKZ results by Dutta et al. [50] and Gao at al. [51]. These authors found that the scaling of the defect density in presence of noise is governed by two independent terms: one noisy term $\sim\tau$, and one KZM term $\sim 1/\sqrt{\tau}$, again different from the logarithm uncovered by our work.
@@ Acknowledgement
DATA AVAILABILITY
The data that support the findings of this article are openly available [72].
Appendix A: Defect generation under noisy ramp fields
As discussed in Sec. 1, defect generation under noisy ramp fields exhibits an AKZ behavior when crossing a quantum critical point, whereby slower driving results in a higher density of defects [50, 51, 52, 53, 54, 55, 56].
@@
As the ramp time scale increases, noise-induced defects dominate the nonadiabatic dynamics, leading to a growth of the defect density with $\tau$. This is indicative of the AKZ regime, where enhancing the ramp time scale results in a higher density of defects in the system. In the limit of a very large values of $\tau$, $n_{\xi}$ is completely governed by the AKZ contribution. In summary, the defect generation is controlled by two competing mechanisms: (1) non-adiabatic defect density, suppressed by increasing the time scale $\tau$ of the ramp; and (2) noise-induced defect density, amplified by increasing $\tau$. As shown in Refs. [50, 51], the interplay of these two competing influences yields a minimum of the density of defects ($n^{\text{min}}_{\xi}$) at an optimum annealing time $\tau_{\text{opt}}$ that scales as $\tau_{\text{opt}}\propto\tau^{-\delta}$, with $\delta=0.666\pm 0.001\simeq 2/3$.
@@
Appendix B: Computing concurrence for time-evolved states of the TFI chain
@@
$C_{\l ,m}(t)=\mbox{max}(0,\lambda_{1}(t)\!-\!\lambda_{2}(t)\!-\!\lambda_{3}(t)\!-\!\lambda_{4}(t)).$ (B1)
@@
$R_{\l ,m}(t)=\sqrt{\sqrt{\rho_{\l ,m}(t)}\tilde{\rho}_{\l ,m}(t)\sqrt{\rho_{\l ,m}(t)}},$ (B2)
@@
Figure 5: (Color online) The logarithm of the density of defects $n_{\xi}$ vs time scale $\tau$ of a ramped quench crossing a quantum critical point in the transverse field Ising chain for different noise intensities $\xi$. The behavior of the density of defects for different values of $\tau$ and $\xi$ illustrates anti-Kibble-Zurek behavior in the presence of a noisy ramp field.
@@
$\rho_{l,m}=\left(\begin{array}[]{cccc}\kappa_{11}&0&0&\kappa_{14}\\ 0&\kappa_{22}&\kappa_{23}&0\\ 0&\kappa_{23}^{\ast}&\kappa_{33}&0\\ \kappa_{14}^{\ast}&0&0&\kappa_{44}\\ \end{array}\right),$ (B7)
@@
$\kappa_{44}$ $=$ $-\langle s^{z}_{l}\rangle+\langle s^{z}_{l}s^{z}_{m}\rangle+\frac{1}{4},$ (B8)
@@
Singling out a reference spin for notational transparency, say at site $\l =1$, and using the JW transformation in Eq. (II), the two-point functions in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) can be expressed in terms of multipoint correlators of the JW fermions,
@@
$\langle s_{1}^{x}s_{1+r}^{y}\rangle$ $\!=\!$ $-\frac{i}{4}\langle B_{1}A_{2}B_{2}...A_{r}B_{r}B_{r+1}\rangle,$ (B9)
@@
$A_{j}=c_{j}^{\dagger}+c_{j},\ \ B_{j}=c_{j}^{\dagger}-c_{j},\ \ j=1,\ldots,r+1.$ (B10)
As for the one-point function $\langle s^{z}_{l}\rangle$ in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain), one has
$\langle s^{z}_{l}\rangle=\frac{1}{2}\langle B_{l}A_{l}\rangle.$ (B11)
The expectation values in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) are defined with respect to the time-evolved fermion states that stand in for the spin states after the JW transformation. For example,
$\langle A_{1}B_{2}\ldots\rangle=\mbox{Tr}(\rho(t)A_{1}B_{2}\ldots)$ (B12)
with $\rho(t)=\otimes_{k}\rho_{k}(t)$, where $\rho_{k}(t)$ is the density operator for mode $k$ in a pure (mixed) ensemble of noise-free (noise-averaged) states from ramped quenches [cf. text before Eqs. (III) and (IV)]. Note that in the case of a pure ensemble, we used the notation $\rho_{k}(t)=\rho_{0,k}(t)$ in Sec. III. Also note that the product over density matrices for single modes runs over $k=\pm(2i-1)\pi/N,i=1,2,\ldots,N/2$ [cf. text after Eq. (2)].
The transverse spin correlation functions in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) can be written on the generic form
$\langle s_{1}^{\alpha}s_{1+r}^{\beta}\rangle=D_{r}^{\alpha\beta}\langle{{\phi_{1}}{\phi_{2}}{\phi_{3}}...\phi_{2r-2}\phi_{2r-1}{\phi_{2r}}}\rangle,$ (B13)
with $D_{r}^{xx}=\frac{1}{4},D_{r}^{yy}=(-1)^{r}\frac{1}{4},D_{r}^{xy}=-\frac{i}{4},$ and $D_{r}^{yx}=(-1)^{r}\frac{i}{4}$, where each operator $\phi_{j},j=1,2,\ldots,2r$, is identified with either $A_{j}$ or $B_{j}$ according to the corresponding expression in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain). Note further that $\langle s_{1}^{z}s_{1+r}^{z}\rangle$ has the simple form
$\langle s_{1}^{z}s_{1+r}^{z}\rangle=\frac{1}{4}\langle{{\phi_{1}}{\phi_{2}}{\phi_{3}}{\phi_{4}}}\rangle,$ (B14)
@@
The expressions in Eqs. (B13) and (B14) are expedient as they allow us to express the multipoint correlators in Eq (Appendix B: Computing concurrence for time-evolved states of the TFI chain) in terms of Pfaffians. This scheme was first exploited by Barouch and McCoy [74] in their study of the XY chain, using a result by Caianiello and Fubini [77] obtained via Wick’s theorem.
@@
where we have written the skew-symmetric matrices on standard abbreviated form. By combining Eqs. (Appendix B: Computing concurrence for time-evolved states of the TFI chain), (B13), (B21), and (B26), replacing the choice l=1l\!=\!1 by the collective index $l$, it is straightforward to show that the correlation functions for nearest-neighbour spins, r=1r\!=\!1, and next-nearest-neighbour spins, r=2r\!=\!2, are given by
@@
$\langle s^{x}_{l}s^{y}_{l+1}\rangle$ $=$ $-\frac{i}{4}\langle B_{l}B_{l+1}\rangle,$ (B27)
@@
$\langle s^{y}_{l}s^{y}_{l+2}\rangle$ $=$ $\frac{1}{4}\large[\langle B_{l+1}A_{l}\rangle\langle B_{l+2}A_{l+1}\rangle\!-\!\langle A_{l}A_{l+1}\rangle\langle B_{l+1}B_{l+2}\rangle-\langle B_{l+2}A_{l}\rangle\langle B_{l+2}A_{l+2}\rangle\large],$ (B28)
@@
$\langle B_{l}B_{\l +r}\rangle$ $=$ $\langle c_{l}^{\dagger}c_{\l +r}^{\dagger}\rangle+\langle c_{l}c_{\l +r}\rangle-\langle c_{l}^{\dagger}c_{\l +r}\rangle-\langle c_{l}c_{\l +r}^{\dagger}\rangle,$ (B29)
@@
where, by way of translational invariance, $r=0,\pm 1,\pm 2$ cover all cases in Eqs. (Appendix B: Computing concurrence for time-evolved states of the TFI chain) and (Appendix B: Computing concurrence for time-evolved states of the TFI chain). We obtain closed expressions for the spin correlations in these equations by calculating the fermion two-point functions in Eq. (Appendix B: Computing concurrence for time-evolved states of the TFI chain). Explicitly, let us consider the two-point function $\langle c_{\l }^{\dagger}c_{\l +r}^{\dagger}\rangle$, suppressing time arguments for notational simplicity, and write
$\langle c_{\l }^{\dagger}c_{\l +r}^{\dagger}\rangle=\mbox{Tr}(\rho\,c_{\l }^{\dagger}c_{\l +r}^{\dagger}),$ (B30)
where, like in Eq. (B12), $\rho=\otimes_{k}\,\rho_{k}$ with $\rho_{k}$ the density operator for mode $k$. By Fourier transforming, one obtains from Eq. (B30),
$\langle c^{\dagger}_{l}c^{\dagger}_{l+r}\rangle$ $=$ $\frac{1}{N}\sum_{p,q}e^{-i(pl+q(l+r))}\langle c^{\dagger}_{p}c^{\dagger}_{q}\rangle$
$=$ $\frac{1}{N}\sum_{p,q}e^{-i(pl+q(l+r))}\mbox{Tr}(\rho\,c^{\dagger}_{p}c^{\dagger}_{q})$